OpenStudy (anonymous):

Which math operation determines the following pattern? 1, 4, 16, 64, ...? addition subtraction multiplication division

OpenStudy (cggurumanjunath):

multiplication.

OpenStudy (cggurumanjunath):

1*4=4 4*4=16 16*4=64...